# Particles before symmetry

### v4

@@ 1 Introduction
Should we value mathematically equivalent formulations of a theory? \@BBOPcitet\@BAP\@BBNFeynman (1994, p. 127)\@BBCP, in his characteristic style, has a persuasive answer to this question:
Every theoretical physicist who is any good knows six or seven different theoretical representations for exactly the same physics. He knows that they are all equivalent, and that nobody is ever going to be able to decide which one is right at that level, but he keeps them in his head, hoping that they will give him different ideas for guessing.
In his Nobel Prize Lecture (“The Development of the Space‑Time View of QED”, 1965), he further reflects on the value of alternative ways of thinking about a theory:
@@
Theories of the known, which are described by different physical ideas may be equivalent in all their predictions and are hence scientifically indistinguishable. However, they are not psychologically identical when trying to move from that base into the unknown. […] If every individual student follows the same current fashion then the variety of hypotheses being generated […] is limited.
The conduit of Feynman’s reflections was his path integral formalism, which was mathematically equivalent to an earlier approach (Schwinger’s) to quantum field theory. The subsequent career of path integrals is evidence of the validity of his arguments.
Another classic case further illustrates the point: Minkowski’s 1908 recasting of Einstein’s special relativity into the language of four-dimensional spacetime geometry. Again, the underlying physics was unchanged, but the shift in formulation was decisive for future developments. What Minkowski introduced was not new predictions but a new ontology: space and time no longer standing apart, but merged into a single structure. And it was precisely this geometrical vantage point that enabled the later generalisation to general relativity (cf. \@BBOPcitep\@BAP\@BBN(Stachel, 2002, p. 226)\@BBCP).
@@
The value of theory-reformulation is a live philosophical topic, admitting a spectrum of positions, from the maximally deflationary instrumentalism to the maximally inflationary fundamentalism (see \@BBOPcitep\@BAP\@BBN(Hunt, 2025)\@BBCP for a recent appraisal). [^1] According to Hunt (p. 5, ibid), instrumentalism takes reformulation to “merely amount to a different choice of convention, no different in kind than a change in notation”; whereas fundamentalism “proposes a metaphysical picture similar to David Lewis’s posits that some properties belong to an elite set of perfectly natural properties, with physics aiming to provide a partial inventory of these”. In this paper, I aim to provide a reformulation that I judge to be valuable under the lights of all but the most deflationary instrumentalist position on this topic.
@@
The reformulation will be of classical gauge theory as it applies to the Standard Model of particle physics, so not quantum electrodynamics or special relativity, and of a less radical nature than either Feynman’s or Minkowski’s reformulations. [^2] This reformulation does not apply to all gauge theories: this will be an important theme in Section 5. Namely, I aim to provide a reformulation of the Standard Model that does not rely on an initial postulate about symmetry; I will call it geometry-first.
@@
The knowledgeable reader will be quick to point out that gauge theory is already highly geometrical: principal fibre bundles and connections are, after all, the stock-in-trade of the geometer. However, the usual formulation also incorporates symmetry into its foundations, and with it, an ontology that extends beyond the spaces where matter fields actually reside.
For theories of particle physics, I will assume the spaces where matter fields reside are vector bundles and that the spaces that introduce symmetry at a ground level are principal fibre bundles. Principal bundles can still appear in a geometry-first formulation, but if they do, since they are not part of the fundamental ontology, they must be entirely supervenient on the structure of the vector bundles, which is the subvenience basis for the geometry-first formulation. I do not demand that the geometry-first formulation should be superior in every respect, or practically advantageous. I will start with a more modest aim: to offer an alternative perspective that clarifies some features of particle theory while omitting symmetry at the ground of the explanatory chain. And, lest I sound too irenic, in Section 5 I will defend my preference for the geometry-first formulation more openly.
@@
Needless to say, symmetry is the cornerstone of particle physics. Representations of Lie groups, Killing forms, Casimir invariants, stabilisers, spontaneous symmetry breaking, gauge-fixing: these are the daily bread of the Standard Model. (This much will be obvious to anyone familiar with the field, so I need not belabor the point.) That the associated principal bundles—and with them the explicit appeal to symmetry at the ground of the explanatory chain—might be dispensed with is therefore anything but trivial.
@@
However, a new geometry-first formulation has recently been proposed, in which symmetries are not postulated and principal fibre bundles are unnecessary \@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025a)\@BBCP. In the alternative formulation, the symmetry groups are only implicit: they arise as the automorphism groups of fundamental vector bundles. The geometry-first formulation is generally available as an alternative only for gauge groups that are linear, and for representations obtained from the fundamental representation (when it is unique) via tensor and direct products, symmetrisation, and other similar operations. Thus the geometry-first formulation demands an alignment between symmetry groups and structure of the vector spaces where matter resides that is not always guaranteed from the symmetry-first perspective, with its largely independent principal fibre bundles and representation spaces. (In Section 5 I will have more to say about the restricted equivalence between geometry-first and symmetry-first formulations).
@@
Even in the cases that admit the two formulations as mathematically equivalent, the geometry-first formulation comes with a significantly different ontology: for the Standard Model of particle physics, it consists of three fundamental vector bundles over spacetime where the various matter fields reside (as sections of tensor products). There is no need for a separate space to encode the principal connections.
@@
Change the formulation, and the explanations change with it. Three examples show how familiar features of particle physics acquire alternative interpretations. First, in a non-Abelian vacuum Yang–Mills theory with Lie group $G$, the fundamental dynamical object is no longer a connection $\varpi$ on a $G$-principal fibre bundle (or its spacetime representative $A_{\mu}^{I}$), but the covariant derivative ${\mathrm{D}}_{\mu}$ on a vector bundle whose automorphism group corresponds to GG—and this remains true even if no vector fields are present to be differentiated: the affine structure of the vector bundle can be dynamical all by itself. In this setting, a particle’s quantum numbers become geometric labels: the internal space it inhabits, and the tensor type it has within that space. (These ideas were explained in \@BBOPcitep\@BAP\@BBN(Gomes, 2024)\@BBCP; so this paper will focus on the next two examples.)
Second, once symmetry groups drop out of the ground level of the explanatory chain, the very notion of ‘symmetry-breaking’ requires reinterpretation. Vector bosons $A_{\mu}^{I}$ are replaced by covariant derivatives of the fundamental bundles, which are not on the same footing as matter fields, and it is no longer clear how they could ‘acquire mass’ in the usual sense. Third, consider the Yukawa couplings. In the symmetry-first formulation, Yukawa terms are scalars formed from sections of different associated bundles, requiring explicit ‘bridges’ between them. In the geometry-first picture, by contrast, the fundamental objects are vector bundles and particle fields are represented by sections of suitable tensor bundles constructed from the fundamental vector bundles. Scalars then arise naturally through inner products and contractions between vectors and their duals.
@@
Another striking example of the explanatory reach of the geometry-first framework is the quantisation of charge. Where the symmetry-first picture attributes it to the topology of the compact group $U(1)$, the geometry-first picture derives it from the discrete algebraic structure of the tensor powers of the fundamental vector bundles. The explanation is at once more general and more direct: geometry enforces discreteness even in settings where symmetry would permit continuity (i.e. for non-compact structure groups).
@@
These examples show how a geometry-first perspective reshapes explanations. Under \@BBOPcite\@BAP\@BBNHunt (2025)\@BBCP’s lights then, this reformulation is non-trivial because it suggests a different ontology, offers new explanations, and displays an epistemic difference in how it solves some problems in common with the usual formulation.333“Successful reformulations clarify what we need to know to solve problems, improving our understanding of the world.” (ibid. p. 5) and “Within [the] shared domain of problems [that they both solve], significant reformulations display an epistemic difference, while trivial reformulations do not.” (ibid. p. 9).
But the real strength of the approach lies in the methodological discipline it suggests. Where the principal–bundle picture allows a loose fit, or slack, between symmetry and geometry, the vector–bundle point of view (VB-POV) ties the two tightly together. Section 5 argues that this apparent restriction is in fact a virtue: it narrows the space of admissible theories in a way that clarifies the ontology of gauge theory and still encompasses our best physics.
Here is how I will proceed. Section 2 introduces both the familiar principal–bundle formulation, which gives a point-of-view on gauge theories (PFB-POV), and the alternative VB-POV. Section 3 gives an alternative gloss of the Higgs mechanism. Section 4 argues that the interpretation of the Yukawa mechanism is more transparent in the geometry-first approach. Section 5 is the most important of this paper: it develops the methodological defense of the VB-POV. Finally, Section 6 draws the broader morals.
@@ 2 Symmetry-first and geometry-first formulations of gauge theory
Here I will give brief overviews of both the familiar, symmetry-first, and of the less familiar, geometry-first formulations of gauge theory. I will start with the more familiar and then introduce the novel.
In the PFB-POV I write the principal connection one-form as $\varpi$; in the VB-POV I write the affine covariant derivative operator as $\nabla$ and the associated-bundle connection as $\omega$.
@@ 2.1 Gauge theory and principal fibre bundles: the symmetry-first formulation
In short, the symmetry-first formulation of the Standard Model is the familiar one: each fundamental interaction is associated with a symmetry group, which is taken as the structure group of a principal fibre bundle. Connections on this bundle then play the role of the vector bosons—the “force carriers.”
Classical configurations of matter particles charged under a force are described by sections of vector bundles associated to the principal bundle whose group encodes that force. One may endow these associated bundles with additional structure (for instance, a Hermitian inner product on $\mathbb{C}^{n}$); in such cases, the representations of the structure group are only required to preserve that structure.
The connection on the principal fibre bundle induces parallel transport on all associated bundles. Crucially, it is the same connection that governs transport in each case, ensuring that different matter fields charged under the same interaction remain coordinated: they all “march in step” under parallel transport, probing the same distributions of electroweak or strong forces. Thus, while associated vector bundles are distinct entities, they are tied together by the principal bundle, which acts as their common coordinator (see \@BBOPcitep\@BAP\@BBN(Weatherall, 2016)\@BBCP and Figure 1). The primacy of the postulated structure group, and the central role it plays, is what makes this a “symmetry-first” formulation.
This is the familiar exposition, but, as we will see, it can be applied more generally. I will say more general symmetry-first formulations of gauge theories in which the symmetries are introduced as part of the principal fibre bundles fall under the principal bundle point of view (PFB-POV).
More technical details are provided in Appendix A, for now, it is sufficient to say that, more generally and in mathematical language, a principal fibre bundle $(P,M,G)$ is a smooth manifold $P$ equipped with a smooth, free action of a Lie group $G$, projecting onto a base manifold MM—spacetime. Intuitively, such a bundle codifies the ways in which the symmetry group $G$ can act on geometric objects defined over $M$. In this paper, I will focus on one especially important class of such objects acted on by the principal bundle: vector bundles. A vector bundle $(E,M,V)$ assigns to each spacetime point $x\in M$ a copy of a fixed vector space $V$, called the typical fibre. Sections of vector bundles are smooth assignments of an element of $V$ to each point of $M$, and matter fields are precisely such sections.
@@
Generally, given a typical fibre $V$ and a principal bundle $(P,M,G)$, we can define associated vector bundles using a representation, i.e. $\rho:G\hookrightarrow GL(V)$ which is determined by $G$, $V$, and the particle lables, or quantum numbers. [^4] Here $\rho$ is understood as an embedding that is, in general, only a partial homomorphism, since it may fail to be faithful or surjective on $\mathsf{Aut}(V)$. This will be important in Section 5. Matter fields are represented as sections of vector bundles associated to the principal bundle whose structure group is the gauge group of the theory (e.g. $SU(3)\times SU(2)\times U(1)$).
The broad idea can be illustrated with the case of a faithful and transitive representation of $G\simeq GL(V)$. We take the points of $P$ to be frames for $V$, understood as ordered tuples—say $V\simeq\mathbb{C}^{n}$—over each point of $M$, so that the group action corresponds to a change of frame on $P$ and a corresponding change of components of a vector in $V$. Since vector fields are frame-independent, one quotients pairs $(p,v)$, with $v\in V$ and $p\in P$, by the simultaneous action of the group on $P$ and on $V$ (via $\rho$). The resulting object is the associated vector bundle,
@@
The advantage of defining associated vector bundles is that it becomes transparent that fields inhabiting bundles associated to the same principal bundle covary under the action of the symmetry group. But it is not the only transparent way of ensuring this covariance, as we will see in Section 2.2.
When introducing the broad idea, it is pedagogical to start with this simple case above, in which, given $E$, the principal fibre bundle $(P,M,G)$ is a frame bundle $L(E)$, with $G\simeq GL(V)$. And we can use the same ideas to get more general principal bundles—call them $L_{\text{\tiny{adm}}}(E)$—when we consider admissible frames for $E$, i.e. frames that respect additional structure on $E$. For instance, if we endow $E$ with a Hermitian metric, we can restrict $L(E)$ to the sub-bundle $L_{\text{\tiny{adm}}}(E)$ of orthonormal frames and recover $(P,M,G)$ with $G\simeq U(V)$, acting in the fundamental representation. We could also endow each fibre with a complex volume element $\varepsilon$, and if $P$ encodes only the symmetry transformations that preserve this coarser geometric structure, we obtain $(P,M,G)$ with $G\simeq SL(V)$—transformations of unit determinant. Alternatively, we could consider transformations preserving both the Hermitian inner product and the volume form, giving $G=SU(V)$, or those rotating only the volume form, giving $G=U(1)$, and so on.
Thus we can form $(P,M,G)$ by picking out subgroups of the group of automorphisms of the typical fiber, $\mathsf{Aut}(V)$. These $(P,M,G)$ are still obtained from the bundle of frames $L(E)$ by restriction to admissible frames.
But the notion of principal and associated vector bundles is not limited to these simple cases. $G$ is not obliged to coincide with $\mathsf{Aut}(V)$; it need not be defined through an associated vector bundle. The definition in (2.1) is valid for any group and representation. Thus, more generally, the structure group of $P$ and the automorphism group of the typical fibre of its associated vector bundles need not be isomorphic—a good thing for my argument, as we will see in Section 5.
For general gauge theories, each associated vector bundle corresponds to a possible particle type, classified by the representation labels that serve as the particle’s quantum numbers. A representation $\rho:G\to GL(V)$ defines an associated vector bundle $E=P\times_{\rho}V$, whose sections $\psi:M\to E$ represent admissible field configurations for a particle type transforming according to $\rho$. The fibre $V$ thus represents the space of internal degrees of freedom (e.g. charge, colour, weak isospin, chirality, etc.), and the representation labels play the role of the particle’s internal “quantum numbers.”
A further, related question is whether such vector bundles merely covary under the group action, or whether they in fact stand in a canonical relation to one another. Answering this will be important to establish a relative advantage of the VB-POV, in later Sections.
Thus, suppose we are given:
@@
In contrast, in the vector-bundle point of view, all the vector bundles that, in the symmetry-first formulation, would be associated to the same principal bundle, are already endowed with a natural relation, as we will now see. Later, in Section 4, we will see how the issue of relating different vector bundles which in the PFB-POV are associated to the same principal bundle can arise in practice.

### v5

@@ 1 Introduction
Mathematically equivalent formulations of a theory can differ profoundly in what they suggest. Feynman made the point in his Nobel Lecture:
@@
Theories of the known, which are described by different physical ideas may be equivalent in all their predictions and are hence scientifically indistinguishable. However, they are not psychologically identical when trying to move from that base into the unknown.
@@
The subsequent career of Feynman’s path integrals—equivalent to Schwinger’s operator methods, yet far more fertile—bears this out. So does Minkowski’s 1908 recasting of special relativity: it changed no predictions, but supplied the geometric ontology without which general relativity could hardly have been conceived.Theory-reformulation is a live philosophical topic, with positions ranging from deflationary instrumentalism to inflationary fundamentalism (\@BBOPcitep\@BAP\@BBN(Hunt, 2025)\@BBCP). [^2] According to Hunt (p. 5, ibid), instrumentalism treats reformulation as “merely a different choice of convention, no different in kind than a change in notation”; fundamentalism “proposes a metaphysical picture similar to David Lewis’s posits that some properties belong to an elite set of perfectly natural properties, with physics aiming to provide a partial inventory of these”. What I offer here should interest anyone who is not a thoroughgoing instrumentalist.
@@
The reformulation I propose applies to classical gauge theory as it bears on the Standard Model—less radical than Feynman’s or Minkowski’s, but nonetheless fruitful, as I will argue. [^3] It does not apply to all gauge theories: this will matter in Section 5. The reformulation does not begin with symmetry so I call it geometry-first.
@@
A knowledgeable reader will protest that gauge theory is already geometrical: principal fibre bundles and connections are, after all, the geometer’s stock-in-trade. True—but the usual formulation also builds symmetry into its foundations, and with it, an ontology that extends beyond the spaces where matter fields actually live.
Here is what I mean. In particle physics, matter fields are sections of vector bundles. The symmetry-first formulation adds principal fibre bundles on top, positing a structure group that coordinates the various matter sectors. I propose to dispense with this extra layer. If principal bundles appear at all in a geometry-first formulation, they must supervene entirely on the vector bundles—not the other way around.
@@
This is not a minor shift. Representations of Lie groups, Killing forms, Casimir invariants, stabilisers, spontaneous symmetry breaking, gauge-fixing: these are the daily bread of the Standard Model. That the explicit appeal to symmetry at the ground of the explanatory chain might be dispensed with is anything but trivial.
@@
But dispensed with it can be. A geometry-first formulation has recently been proposed in which symmetries are not postulated and principal fibre bundles are unnecessary (\@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025b)\@BBCP). Symmetry groups become implicit: they arise as automorphism groups of fundamental vector bundles. The formulation is available only for gauge groups that are linear, and for representations built from the fundamental representation via tensor products, symmetrisation, and the like. It therefore demands an alignment between symmetry and geometry that the symmetry-first picture does not guarantee.
@@
Even when the two formulations are mathematically equivalent, they come with different ontologies. For the Standard Model, the geometry-first picture posits three fundamental vector bundles over spacetime. Matter fields are sections of tensor products of these bundles. There is no separate space encoding the principal connection.
Change the formulation, and the explanations change with it. Four examples, developed below, show how.
@@
First, in a non-Abelian vacuum Yang–Mills theory with Lie group $G$, the fundamental dynamical object is no longer a connection $\varpi$ on a $G$-principal bundle (or its spacetime representative $A_{\mu}^{I}$), but the covariant derivative $\nabla_{\mu}$ on a vector bundle whose automorphism group is $G$. This remains true even when no matter fields are present: the affine structure can be dynamical all by itself. Quantum numbers become geometric labels—the internal space a particle inhabits, and its tensor type within that space. (These ideas were explained in \@BBOPcitep\@BAP\@BBN(Gomes, 2024)\@BBCP; this paper focuses on the next three.)
Second, once symmetry groups drop out of the ground level, ‘symmetry-breaking’ requires reinterpretation. Vector bosons are replaced by covariant derivatives of fundamental bundles, which are not on the same footing as matter fields. How could they ‘acquire mass’ in the usual sense?
Third, Yukawa couplings. In the symmetry-first formulation, Yukawa terms are scalars formed from sections of different associated bundles, requiring explicit bridges between representation spaces—bridges that need not be unique. In the geometric picture, these scalars arise through inner products and contractions that are already determined by the bundle structures.
@@
Fourth, the quantisation of charge. The symmetry-first picture attributes it to the topology of the compact group $U(1)$. The geometric picture derives it from the discrete algebraic structure of tensor powers of the fundamental bundles. This means that even for non-compact groups, where symmetry would permit continuity, geometry enforces discreteness.
@@
Under \@BBOPcitet\@BAP\@BBNHunt (2025)\@BBCP’s lights, such a reformulation is non-trivial: it suggests a different ontology, offers new explanations, and displays an epistemic difference in how it solves problems.444“Successful reformulations clarify what we need to know to solve problems, improving our understanding of the world” (ibid. p. 5); “Within [the] shared domain of problems [that they both solve], significant reformulations display an epistemic difference, while trivial reformulations do not” (ibid. p. 9).
A word of caution about scope. The VB-POV applies to a proper subclass of gauge theories: those whose gauge groups arise as automorphism groups of structured vector spaces, with matter built tensorially from fundamental bundles. As a theoretical framework, this is genuinely narrower than the PFB-POV—it excludes models that the principal-bundle formulation can perfectly well accommodate. But the Standard Model lies within the overlap, and within that overlap the two formulations are empirically equivalent. What changes is which features of the physics appear primitive and which derived. Two kinds of claim should therefore be kept apart throughout: claims about explanatory structure—the Higgs and Yukawa reformulations of Sections 34—and claims about admissible theory-space—notably the slack diagnosis and the charge-quantisation discussion of Section 5. The former are reformulations in the strict sense; the latter amount to extra constraints that I argue physicists already tacitly impose.
Section 2 reviews both formulations. Sections 3 and 4 rework the Higgs mechanism and the Yukawa couplings from the geometry-first standpoint. Section 5 argues that the apparent narrowness of the geometry-first picture is its chief virtue. Section 6 draws lessons.
@@ 2 Symmetry-first and geometry-first formulations of gauge theory
I begin with the familiar symmetry-first formulation, then introduce the geometry-first alternative. A word on notation: $\varpi$ always denotes the principal connection one-form on $P$; $\nabla$ denotes a covariant derivative on a vector bundle $E$; and $\omega$ denotes the local connection form on $E$ that appears upon choosing a frame, so that $\nabla={\mathrm{d}}+\omega$ with $\omega\in\Gamma(T^{*}M\otimes\mathsf{End}(E))$. When I speak of the affine structure of a vector bundle, I mean the space of covariant derivatives compatible with the fibre structure (e.g. Hermitian metric): this space is an affine space modelled on $\Omega^{1}(\mathsf{End}(E))$, so that any two compatible covariant derivatives differ by an endomorphism-valued one-form.
@@ 2.1 Gauge theory and principal fibre bundles: the symmetry-first formulation
The symmetry-first formulation of the Standard Model is the familiar one. Each fundamental interaction is associated with a symmetry group, taken as the structure group of a principal fibre bundle. Connections on this bundle play the role of vector bosons—the force carriers.
Classical configurations of matter particles charged under a force are sections of vector bundles associated to the principal bundle whose group encodes that force. One may endow these associated bundles with additional structure—a Hermitian inner product on $\mathbb{C}^{n}$, say—in which case the representations of the structure group need only preserve that structure.
The connection on the principal bundle induces parallel transport on all associated bundles. Crucially, it is the same connection in each case: different matter fields charged under the same interaction march in step under parallel transport, probing the same distribution of electroweak or strong forces. Associated vector bundles are distinct entities, but the principal bundle ties them together, acting as their common coordinator (see \@BBOPcitep\@BAP\@BBN(Weatherall, 2016)\@BBCP and Figure 1). The primacy of the postulated structure group is what makes this symmetry-first.
Any formulation of gauge theory that introduces symmetries via principal bundles falls under what I call the principal bundle point of view (PFB-POV). Technical details are in Appendix A. For now: a principal fibre bundle $(P,M,G)$ is a smooth manifold $P$ equipped with a smooth, free action of a Lie group $G$, projecting onto a base manifold $M$ (spacetime). Such a bundle codifies the ways $G$ can act on geometric objects over $M$. The most important such objects for this paper are vector bundles. A vector bundle $(E,M,V)$ assigns to each spacetime point $x\in M$ a copy of a fixed vector space VV—the typical fibre. Matter fields are sections: smooth assignments of an element of $V$ to each point of $M$.
@@
Generally, given a principal bundle $(P,M,G)$ and a representation $\rho:G\to GL(V)$ on a typical fibre $V$, one constructs an associated vector bundle $E=P\times_{\rho}V$ (with $\rho$ not necessarily faithful or surjective onto $\mathsf{Aut}(V)$ [^5] Here $\rho$ is an embedding that may fail to be faithful or surjective onto $\mathsf{Aut}(V)$.). In this framework, each such bundle corresponds to a particle type, its sections $\psi:M\to E$ are admissible field configurations, and the representation labels encode internal quantum numbers such as charge, colour, and weak isospin; for the Standard Model, the relevant gauge group is $SU(3)\times SU(2)\times U(1)$.
The resulting object is the associated vector bundle,
@@
The advantage of this construction: fields in bundles associated to the same principal bundle manifestly covary under the symmetry group. (But, as we will see in Section 2.2, it is not the only route to covariance).
The simplest case is $G\simeq GL(V)$, where $P$ is the full frame bundle of a vector bundle $E$. Enriching the fibre with further structure—a Hermitian metric, a complex volume form, or both—restricts the admissible frames and thereby selects a subgroup of $GL(V)$ as the structure group (see Appendix A for details).
But the notion of principal and associated vector bundles is not limited to these simple cases. $G$ is not obliged to coincide with $\mathsf{Aut}(V)$; it need not be defined through an associated vector bundle. The definition in (2.1) is valid for any group and representation. The structure group of $P$ and the automorphism group of the typical fibre need not be isomorphic. This will be important in Section 5.1.
A further question: do such vector bundles merely covary under the group action, or do they stand in canonical relation to one another? Suppose we are given:
@@
In contrast, in the geometry-first picture, all vector bundles that in the symmetry-first formulation would be associated to the same principal bundle are already endowed with natural relations, as we now see.